OpenStudy (anonymous):

factor completely. factor out the GCF first. then apply either the difference of 2 squares method or trial and error 1.2n^2+18n+28 2.3x^2-12x-15 3.5x^2-20x+20 4.2x^2-30x+100 5.3x^2-21x+30 6.10n^2+60n+80 7.4x^2+24x+32 8.2x^2+2x-4 9.x^3+7x^2+6x 10.5x^2-20 11.x^3-20 12.x^4-9x^2 13.y^4-4y^2 14.x^3-49x

OpenStudy (anonymous):

#1 2n^2+18n+28 =0 2 * (n^2 + 9n +14) = 0 2* (n+2) * (n+7) =0 =====> n= -2 or n= -7

OpenStudy (anonymous):

#2. 3x^2-12x-15 =0 3 * ( x^2 -4x -5 ) =0 3 * (x+1) * (x-4) = 0 ====> x =-1 or x =4

OpenStudy (anonymous):

#3. 5x^2-20x+20 =0 5 * ( x^2 -4x +4) =0 5*( x-2)*(x-2) =0 =====> x =2